A 1.55 kg rock is suspended from a spring that is stretched 0.870 m. What is the spring constant for this spring?



1
Expert's answer
2020-12-31T08:08:41-0500

By definition, the spring constant is:


"k=\dfrac{P} {x}"

where "P=mg" is the weight of the rock, "x=0.87m" is the elongation of the string, "m=1.55kg" is the mass of the rock, and "g=9.81m\/s^2" is the gravitational acceleration. Thus, obtain :


"k=\dfrac{mg} {x} \\
k=\dfrac{1.55\cdot9.81} {0.87}\approx17.48N/m"

Answer. 17.48 N/m.